OpenStudy (anonymous):

Solve the equation 2x+1 = 30. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

Ok, to solve, you want to find the value of x.

OpenStudy (anonymous):

so for the first step you want to move the +1 to the right side, 30-1=29, then you want to cancel the 2 on the left side, you bring it to the right side which makes it 29/2, so your answer is x=29/2